OpenStudy (anonymous):

Write the equation of a line with slope m=1/2 and including point (4, -1).

OpenStudy (anonymous):

plug those things given to you into point slope form: \(\large y-y_0=m(x-x_0) \) where m is the given slope, and (x0, y0) is the given point.

OpenStudy (anonymous):

y = mx + b where m = slope b = y-inercept

OpenStudy (anonymous):

B. y = 2x + 3 ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

to find b, plug in (4, -1) in to y = mx + b